OpenStudy (anonymous):

i need help with this problem.. subtract (8+7i)-(4-6i)????

OpenStudy (anonymous):

Their complex numbers. Do you know what that is ?

OpenStudy (anonymous):

This is the problem of complex number.A complex number has real and imaginary part so to solve this put together the real and imaginary parts (8+7i)-(4-6i)=8-4+7i+6i=4+13i

OpenStudy (anonymous):

@ dpasigh thanks for helping me out